OpenStudy (anonymous):

6 over 4=x. whats the answer in decimal form though?

OpenStudy (anonymous):

@Opcode

OpenStudy (anonymous):

1.5

OpenStudy (zzr0ck3r):

if you break 1 dollar into 4 what do you have?

OpenStudy (opcode):

\(\huge \frac {3} {2} = x\) \(1.5\)

OpenStudy (zzr0ck3r):

if you have 6 of those fourths what do you have?

OpenStudy (anonymous):

Its 1.5 .

OpenStudy (zzr0ck3r):

yes but you should lead them to the answer not just give it to them, otherwise we are not helping:)

OpenStudy (anonymous):

thanks but yes i do need help with figuring it out though lol

OpenStudy (zzr0ck3r):

how much is 6 quarters?

OpenStudy (zzr0ck3r):

in dollars

OpenStudy (anonymous):

1.25$

OpenStudy (zzr0ck3r):

thats 5

OpenStudy (zzr0ck3r):

how much is 6?

OpenStudy (zzr0ck3r):

you just told me 5/4 = 1.25 and note how 6 quarters is 1.50 in other words...6/4 = 1.5

OpenStudy (opcode):

@zzr0ck3r Wouldn't it be easier to just show the steps?: \(\huge \frac {6} {4} = x\) This simplifies to: \(\huge \frac {3} {2} = x\) \(\huge \frac {3} {2} = 1.5\) \(\huge x = 1.5\) This way you don't get confused with all the money and stuff.

OpenStudy (anonymous):

okay i understand. i dont know this one though. 11.5 over 4.

OpenStudy (zzr0ck3r):

um if you dont know 6/4 = 1.5 i doupt you know 3/2 is 1.5 ....

OpenStudy (zzr0ck3r):

for that you would want to do long division

OpenStudy (zzr0ck3r):

@Opcode I find students lern the metric best by using money

OpenStudy (opcode):

Well that always used to confuse me :P I guess people learn differently :)

OpenStudy (zzr0ck3r):

are you in the usa?

OpenStudy (zzr0ck3r):

or any other place that uses metric for money?

OpenStudy (anonymous):

Im in the usa.